OpenStudy (anonymous):

Find the equation of a circle with centre (0,0) that passes through (8,-15)

OpenStudy (anonymous):

x^2 + y ^2 = r ^2 (substitute the values :D

OpenStudy (anonymous):

can yu show me plug in the values

OpenStudy (anonymous):

\[x^2 + y ^2 = r ^2\] the length of the radius of the circle is \[r=\sqrt{(8)^2+(-15)^2} = 17\] \[x^2+y^2=289\] <-- is the equation of the circle

OpenStudy (anonymous):

lol fine -.- (8)^2 + (-15)^2 = r^2 64 + 225 = r ^2 289 = r^2 13 = r

OpenStudy (anonymous):

where did yu get 17 from

OpenStudy (anonymous):

its 13 not 17

OpenStudy (anonymous):

nvm its 17

OpenStudy (anonymous):

how do yu kno

OpenStudy (anonymous):

lol kay hes right its 17

OpenStudy (anonymous):

he took the square root of 17

OpenStudy (anonymous):

i mean 289

OpenStudy (anonymous):

yeh, just look at the answer up there and you will know where the 17 came from, it is the magnitued of the radius, whenever you want to find the magnited of a vector lets say (2,3), it is equal to the square root of the x, y values square. take a look at this: http://www.mathsisfun.com/algebra/circle-equations.html

OpenStudy (anonymous):

), it is equal to the square root of the x, y values SQUARED**